OpenStudy (anonymous):

3(sqrt)((2z)^30) simplify?

OpenStudy (farmdawgnation):

I'm having trouble understanding the original equation.... do you mean: \[3\sqrt{2z ^{30}}\] or something else?

OpenStudy (akshay_budhkar):

yes or is it a cube root?

OpenStudy (anonymous):

Yes, this is a cube root.

OpenStudy (anonymous):

i am going to make a bet that it is \[\sqrt[3]{(2z)^{30}}\]

OpenStudy (farmdawgnation):

How far did you get Leibymisty?

OpenStudy (anonymous):

yes thanks I don't know how to put in cube root, thanks lol

OpenStudy (anonymous):

That's where my be is too: \[\sqrt[3]{(2z)^{30}}\]

OpenStudy (anonymous):

in which case think " thirty divided by 3 is 10" and write \[(2z)^{10}\]

OpenStudy (anonymous):

Im doing my homework on a website and I've tried to enter (2z)^10 and it's telling me its incorrect...

OpenStudy (anonymous):

2^10 = 1024 so maybe 1024z^10

OpenStudy (farmdawgnation):

@Leibymisty - if your website is anything like the system I used in college, double check that you can use two digits in an exponent without parenthesis. It might require you type: (2z)^(10) Backwards, I know, but it happens.

OpenStudy (anonymous):

TBates you're a genius...Thanks! and thanks to all the rest who helped!

OpenStudy (farmdawgnation):

Anytime. :)

OpenStudy (anonymous):

Well naturally, genius is what I do best. =P